OpenStudy (anonymous):

Which of the following is a polynomial with roots – square root of 3, square root of 3, and –2 ? x^3 – 2x^2 – 3x + 6 x^3 – 3x^2 – 5x + 15 x^3 + 2x^2 – 3x – 6 x^3 + 3x^2 – 5x – 15

OpenStudy (solomonzelman):

\[roots:~~±\sqrt{3},~2\] Have you tried to solve the problem? I would like to see your work on this.

OpenStudy (solomonzelman):

-2, not just 2... anyway, .. what have you attempted ?

OpenStudy (anonymous):

I gave the options but I need this explained to me, sorry. Do you know how to walk me through it ?

OpenStudy (anonymous):

Or i mean can you ?

OpenStudy (solomonzelman):

The first step would be to says \[f(x)=(x^2-3)(x+2)\]because \[x^2-3=0\] gives that \[x= ±\sqrt{3}\]the 2 of our roots. then \[x+2=0\]gives x = -2 So therefore we say that the function is \[f(x)=(x^2+3)(x+2)\]You just need to finish this, by expending the right hand side. Sorry for not helping -:(

OpenStudy (anonymous):

Oh! Nevermind, I figured it out! Thanks, though! (put the numbers in place and see what equation equates to zero!)